Please help me solve the following problem in detail. I am trying to help my daughter and I am the worst at algebra.
6x-4(3x-2)=10

Answer by longjonsilver(2297)

6x-4(3x-2)=10
6x-12x+8 = 10
-6x+8 = 10
-6x = 2
x = 2/(-6)
--> x = -1/3

jon.